On January 1, 2018, the Excel Delivery Company purchased a delivery van for $33,000. At the end of its five-year service life, it estimated that the van will be worth $3,000. During the five-year period, the company expects to drive the van 100,000 miles Required Calculate annual depreciation for the five-year life of the van using each of the following methods. 1. Straight line 2. Sum-of-the-years-digits. 3. Double-declining balancee 4. Units of production using miles driven as a measure of output, and the following actual mileage Complete this question by entering your answers in the tabs below Required Required 2 Required 3 Required 4 Calculate annual depreciation for the five-year life of the van using units of production using miles driven as a measure of output, and the following actual mileage. (Do not round your intermediate calculations.) Year 2018 2019 2020 2021 2022 Total Miles Depreciatio 22,000 24,000 15,000 20,000 21,000
Answer #1
1
Straight line depreciation per year = (33000-3000)/5= $6000
2
Sum-of-the-years-digits = 5*(5+1)/2= 15
Year Depreciation
2018 10000 =(33000-3000)/15*5
2019 8000 =(33000-3000)/15*4
2020 6000 =(33000-3000)/15*3
2021 4000 =(33000-3000)/15*2
2022 2000 =(33000-3000)/15*1
3
Double declining rate =(1/5)*2 = 40%
Year Beginning book value Rate Depreciation expense Accumulated depreciation Ending book value
2018 33000 40% 13200 13200 19800
2019 19800 40% 7920 21120 11880
2020 11880 40% 4752 25872 7128
2021 7128 40% 2851 28723 4277
2022 4277 40% 1277 30000 3000
4
Depreciation rate per mile = (33000-3000)/100000= $0.30
Year Miles Depreciation
2018 22000 6600 (22000*0.30)
2019 24000 7200 (24000*0.30)
2020 15000 4500 (15000*0.30)
2021 20000 6000 (20000*0.30)
2022 21000 5700 =30000-(6600+7200+4500+6000)
Total 30000